# Dynamic Input Mapping Inversion for Algebraic Loop-Free Control in Hydraulic Actuators

### v2

@@ III-C2 Control input derivation
Directly isolating $u$ in Eq. 11 and Eq. 13 is not possible: while it may be possible to invert the $K(u)$ function in Eq. 3, the control input also appears inside the Heaviside functions, generating an algebraic loop where the derived input depends on the input itself. The standard way to solve the problem is the one described in [4]:
$Q_{p}(t,u)=$ $K(u)\left[\sqrt{p_{s}-p_{p}}\,H\left(u\right)+\sqrt{p_{p}-p_{t}}\,H\left(-u\right)\right]$ (17)
$\approx$ $Ku\left[\sqrt{p_{s}-p_{p}}\,H\left(u_{-1}\right)+\sqrt{p_{p}-p_{t}}\,H\left(-u_{-1}\right)\right]$
@@
$Q_{r}(t,u)=$ $K(u)\left[\sqrt{p_{r}-p_{t}}\,H\left(u\right)+\sqrt{p_{s}-p_{r}}\,H\left(-u\right)\right]$ (18)
$\approx$ $Ku\left[\sqrt{p_{r}-p_{t}}\,H\left(u_{-1}\right)+\sqrt{p_{s}-p_{r}}\,H\left(-u_{-1}\right)\right]$
@@
the standard feedback linearization controller is then,
$u_{FL}=\frac{1}{g}\left[\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}\right],$ (20)
and the standard super-twisting sliding mode controller is,
$u_{ST}=\frac{1}{g}$ $\left[\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}\right.$ (21)
$\,\,\left.+r^{(3)}-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s)d\tau\right].$
The $K(u)$ linearization is then reflected into uncertainties in the input term $g$ and in the terms $K_{Qp}$ and $K_{Qr}$.
@@ IV Cascaded control design
The delay operator introduced to avoid the formation of algebraic loops simplified the controller implementation with the drawback of increasing unwanted chatter on the control signal, with the consequences on oil quality described above. The solution proposed in this article for avoiding this, is to leverage a virtual input for the controller design and dynamically deriving a real input that produces the virtual one required by the controller. The block diagram is shown in Fig. 2.
Figure 2: Block diagram of the proposed algorithm: Cascaded control with dynamic input mapping inversion.
By defining a virtual input as
@@
where the virtual input $v$ is a function of time and the control input $u$, and is now present as a linear term. The feedback linearizing and supertwisting control designs are now defined in terms of the associated virtual inputs as, respectively,
$v_{FL}=\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}},$ (24)
and
$v_{ST}=$ $\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}$ (25)
$+r^{(3)}-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))\,d\tau.$
@@ IV-A Dynamic mapping inversion algorithm
The last part of the proposed method constitutes the derivation of the control input $u$ by solving Eq. 22 for $u$ in each case. A dynamic nonlinear mapping inversion scheme is adopted from [16] for this purpose, allowing for re-formulating of the input calculation into a parameter estimation problem in $u$ for a nonlinearly parametrized regressor.
@@
be a mapping from any given actual control input $u$ to the virtual input that was earlier defined in Eq. 22. Assume that for every value $v$ at each time instant there exists an appropriate value of the control input $u^{*}$ such that $v=\rho(u^{*})$. It has been shown [16, 17] that the estimation law
@@
If the mapping $\rho$ is not injective, then under Conditions 1-3, the estimation law in Eq. 27 will ensure that $u$ converges to an appropriate value such that $\rho(t,u)$ converges to $v$ exponentially fast.
Figure 3: Lookup table approximation with polynomial functions.
A typical choice for $\mu$ is the gradient of $\rho$ with respect to $u$. However, computation of $\mu$ requires differentiation of the nonlinear function $K(u)$ present in Eq. 3 and, since the latter is usually available as a lookup table, it is approximated by an appropriate polynomial function of $u$. Fig. 4 illustrates an example of different polynomial approximations for $K(u)$, i.e.
$K(u)=\sum_{i=0}^{n}\alpha_{i}\,u^{i},$ (31)
with a fifth-order polynomial yielding the best fit.
This particular approximation also ensures that $\frac{\partial K}{\partial u}>~{}0,\;\forall u\in[-1,1]$ as shown in Fig. 4, which will be useful later in proving the validity of the convergence conditions. The function of $\mu$ can now be selected as
@@
Moreover, since $\frac{\partial\rho}{\partial u}$ is bounded, there exists a Lipschitz constant $L_{u}>0$ such that
@@
Figure 4: Verification of derivative for algorithm convergence condition.
@@ V Stability Analysis
The stability of the closed-loop system is analyzed by including the nonlinear controller, the dynamical mapping inversion, and the disturbance estimator. By writing it in the form of a feedback interconnected system
@@
and by fulfilling the following conditions:
Condition 4.
Unperturbed $\Sigma_{1}$ $(\tilde{u}=0)$ is Uniformly GES
Condition 5.
Solutions of $\Sigma_{2}$ are Uniformly Ultimately Bounded (UUB)
Condition 6.
$\forall(t,\boldsymbol{\zeta})\in[0,\infty)\times\mathbf{R}^{2}$
$\frac{\partial V(t,\boldsymbol{\zeta})}{\partial\boldsymbol{\zeta}}h(t,\tilde{u})=o(W(\boldsymbol{\zeta})),$ (36)
the overall closed-loop system is globally UUB according to Theorem 1 from [21].
Let $v^{*}=\rho(u^{*})$ be the required virtual controller and $u^{*}$ the related control input, one can define
$\tilde{v}\triangleq v^{*}-v=\rho(u^{*})-\rho(u),$ (37)
for some arbitrary $u$ and $v$, with $\tilde{u}=u^{*}-u$ having been defined earlier. The two closed-loop systems obtained by closing the loop with the FL controller and the ST-SMC controller are, respectively:
$\Sigma_{1,FL}:\dot{\underline{\mathbf{e}}}=(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{\xi_{F_{d}}}+\boldsymbol{h_{FL}}(t,\tilde{u})$ (38)
$\Sigma_{2,FL}:\dot{\tilde{u}}=\dot{u}^{*}-\gamma\,\mu(x,u^{*},\tilde{u})\,\Big{(}-K_{FL}\underline{\mathbf{e}}-\rho(u^{*}-\tilde{u})\Big{)}$
@@
$\boldsymbol{\xi}_{F_{d}}=\boldsymbol{B_{FL}}\left(-\frac{\dot{\tilde{F}}_{d}}{M}+K_{FL}\begin{bmatrix}0\\ -\displaystyle\frac{\tilde{F}_{d}}{M}\end{bmatrix}\right)$ (39)
$\boldsymbol{h_{FL}}(t,\tilde{u})=-\boldsymbol{B_{FL}}\,\rho(\tilde{u}),$
and
$\Sigma_{1,ST}:\dot{\boldsymbol{\vartheta}}=f_{\boldsymbol{\vartheta}}(t,\boldsymbol{\vartheta})+\boldsymbol{h_{ST}}(t,\tilde{u})$ (40)
$\Sigma_{2,ST}:\dot{\tilde{u}}=\dot{u}^{*}-\gamma\,\mu(x,u^{*},\tilde{u})\,\Big{(}v^{*}(t,\boldsymbol{\vartheta})-\rho(u^{*}-\tilde{u})\Big{)}$
with
$f_{\boldsymbol{\vartheta}}(t,\boldsymbol{\vartheta})$ $=\frac{\partial T}{\partial\boldsymbol{\zeta}}\left\{\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\frac{k_{1}}{2}&1\\ -k_{2}&0\end{bmatrix}T^{-1}(\boldsymbol{\vartheta})-\begin{bmatrix}0\\ d\end{bmatrix}\right\}$ (41)
$\boldsymbol{h_{ST}}(t,\tilde{u})$ $=-\frac{\partial T}{\partial\boldsymbol{\zeta}}\begin{bmatrix}\displaystyle\frac{1}{|\zeta_{1}|}\\ 0\end{bmatrix}\rho(\tilde{u}-\hat{u})$
according to [22], where $v^{*}(t,\boldsymbol{\zeta})$ was defined in Eq. 25 and $\boldsymbol{\zeta}$ corresponds to a coordinates transformation of the closed-loop dynamics. The detailed derivations are presented in Appendix.
Condition 4 coincides with the proof of stability of the controllers in their standard configuration, and is ensured by choosing the Lyapunov functions as:
$V_{FL}$ $=\underline{\mathbf{e}}^{T}P_{FL}\,\underline{\mathbf{e}}$ (42)
$V_{ST}$ $=\boldsymbol{\vartheta}^{T}P_{ST}\,\boldsymbol{\vartheta}.$
Condition 5 is satisfied according to the result in [23], where it was proven that if $|\dot{u}^{*}|<\infty\,\forall t>0$, then $\tilde{u}(t)$ is UUB for all time.
Finally, Condition 6 is verified as, $\Sigma_{1,FL}$:
@@
$\left\|\frac{\partial V_{FL}}{\partial\underline{\mathbf{e}}}\boldsymbol{h_{FL}}(t,\tilde{u})\right\|$ $\leq\left\|2P_{FL}\underline{\mathbf{e}}\right\|\,\left\|B_{FL}\right\|\,\left|\rho(u^{*})-\rho(u^{*}-\tilde{u})\right|$ (43)
$\leq\left\|2P_{FL}\underline{\mathbf{e}}\right\|\,\left\|B_{FL}\right\|\,L_{p}\,\left\|\tilde{u}\right\|$
@@
$\leq\underbrace{\left\|2P_{FL}\right\|\,\left\|B_{FL}\right\|\,L_{p}\,\left\|\tilde{u}\right\|_{\infty}}_{\eta}\,\left\|\underline{\mathbf{e}}\right\|,$
and for $\Sigma_{1,ST}$,
$\left\|\frac{\partial V_{ST}}{\partial\boldsymbol{\vartheta}}\boldsymbol{h_{ST}}(t,\tilde{u})\right\|$ $\leq\left\|2P_{ST}\boldsymbol{\vartheta}\right\|\frac{L_{\rho}}{\min_{t}(|\zeta_{1}(t)|)}\|\tilde{u}\|_{\infty}\left\|\frac{\partial\tau}{\partial\boldsymbol{\xi}}\right\|_{\infty}$ (44)
$\leq\underbrace{\left\|2P_{ST}\right\|\frac{L_{\rho}}{\min_{t}(|\zeta_{1}(t)|)}\|\tilde{u}\|_{\infty}\left\|\frac{\partial\tau}{\partial\boldsymbol{\xi}}\right\|_{\infty}}_{\eta^{\prime}}\left\|\boldsymbol{\vartheta}\right\|.$
@@ VI Case Study
$J_{tot}\dot{v_{\beta}}=\left(A_{p}p_{p}-A_{r}p_{r}\right)r\sin\left(\psi\left(x_{c}\right)\right)-M_{fric}-M_{ext}$ (45)
@@
where the equivalent mass $M$ also accounts for blade inertia. It is noticed that the equations match the dynamics of the cylinder dynamics presented in Eq. 7. The the presence of a cyclic component in the variable distance that relates mechanical moment to force is minor, as will be shown. Thus, the control design aims at tracking the angle while the cylinder dynamics represents a stable internal dynamics, under the assumption of having a stiff mechanical connection between cylinder and blade. Hence, there exists a static transformation relating $x_{c}$ to $\beta$ and with the pitch control error being the difference between pitch angle and a reference,
$e=\beta-r,$ (46)
the relative degree of the closed loop pitch control will be the same as for the closed loop cylinder piston position control loop analysed above. With relative degree three of both, the third derivative of the error is needed,
$e^{(3)}=$ $-\frac{\dot{M}_{fric}+\dot{M}_{ext}}{J_{tot}}$ (47)
$-\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}\sin{\psi_{p}(x_{c})}-\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}\sin{\psi_{p}(x_{c})}$
$+\frac{lA_{p}B}{J_{tot}V_{p}}Q_{p}(u)\sin{\psi_{p}(x_{c})}+\frac{lA_{r}B}{J_{tot}V_{r}}Q_{r}(u)\sin{\psi_{p}(x_{c})}$
$+\underbrace{\frac{l}{J_{tot}}F_{h}\cos(\psi_{p}(x_{c}))\frac{d\psi}{dx_{c}}v_{c}}_{\tau(x)}-r^{(3)}\,\triangleq\,\varpi.$ (48)
With respect to the general case presented in Eq. 8, two variations are present:
  • $1/M$ is replaced by $l\sin{\psi_{p}(x_{c})}/J_{tot}$.
@@
However, Fig. 5 suggests that $\psi(x_{c})$ is approximately constant, and thus the following assumptions are made:
  • $\psi(x_{c})\approx\textit{const}$
  • $\frac{d\psi}{dx_{c}}\approx 0\rightarrow\tau(x)\approx 0$.
VI-A Standard controllers for study case

### v4

@@ III-A Standard control input derivation
Directly isolating $u$ in (11) is not possible: while it may be possible to invert the $K(u)$ function in (3), the control input also appears inside the Heaviside functions, generating an algebraic loop where the derived input depends on the input itself. The standard way to solve this problem was described by [ayalewCascadeTuningNonlinear2006],:
$Q_{p}(t,u)=$ $K(u)\left[\sqrt{(p_{s}-p_{p})\,H(u)}+\sqrt{(p_{p}-p_{t})\,H(-u)}\right]$ (12)
$\approx$ $Ku\left[\sqrt{(p_{s}-p_{p})\,H(u_{-1})}+\sqrt{(p_{p}-p_{t})\,H(-u_{-1})}\right]$
@@
$Q_{r}(t,u)=$ $K(u)\left[\sqrt{(p_{r}-p_{t})\,H(u)}+\sqrt{(p_{s}-p_{r})\,H(-u)}\right]$ (13)
$\approx$ $Ku\left[\sqrt{(p_{r}-p_{t})\,H(u_{-1})}+\sqrt{(p_{s}-p_{r})\,H(-u_{-1})}\right]$
@@
a standard feedback linearizing control is obtained by,
$u=\frac{1}{g}\left[\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}\right].$ (15)
The $K(u)$ linearization is then reflected into uncertainties in the input term $g$ and in the terms $K_{Qp}$ and $K_{Qr}$ from (12) and (13).
@@ IV Cascaded control design
The delay operator introduced to avoid the formation of algebraic loops simplified the controller implementation but have the drawback of increasing unwanted chatter on the control signal [ayalewCascadeTuningNonlinear2006],with associated consequences on oil quality degradation that were mentioned above. The solution proposed in this article, to avoid this, is to leverage a virtual input for the controller design and dynamically deriving a real input that produces the virtual one required by the controller. The block diagram for the proposed algorithm is shown in Fig. 3.
Defining a virtual input
@@
where $v$ is a function of time and the control input $u$, and is now present as a linear term. A feedback linearizing controller is now defined in terms of an associated virtual input as
$v=\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}.$ (18)
@@ IV-A Dynamic mapping inversion algorithm
The last part of the proposed method constitutes the derivation of the control input by solving (16) for $u$. A dynamic nonlinear mapping inversion scheme is adopted from [gripParameterEstimationCompensation2010b] for this purpose, allowing for re-formulating of the input calculation into a parameter estimation problem in $u$ for a nonlinearly parametrized regressor.
@@
Figure 3: Block diagram of the proposed algorithm: Cascaded control with dynamic input mapping inversion. The input inversion is explained below.
be a mapping from any given actual control input $u$ to the virtual input that was defined in (16). Assume that for every value $v$ at each time instant there exists an appropriate value of the control input $u^{*}$ such that $v=\rho(u^{*})$. It has been shown [gripParameterEstimationCompensation2010b, papageorgiouRobustBacklashEstimation2019] that the estimation law
@@
Figure 4: Manufacturer’s lookup table of valve characteristic and approximation with a polynomial function.
The final control action is then given by
$u(t)=\gamma\int_{0}^{t}\mu\big(\boldsymbol{x}(\tau),u(\tau)\big)\,\Big(v(\tau)-\rho\big(\boldsymbol{x}(\tau),u(\tau)\big)\Big)d\tau,$ (24)
where $v$ is the one defined in (18).
@@
If the mapping $\rho$ is not injective, then under Conditions 1-3, the estimation law in (20) will ensure that $u$ converges to an appropriate value such that $\rho(t,u)$ converges to $v$ exponentially fast.
A typical choice for $\mu$ is the gradient of $\rho$ with respect to $u$. However, computation of $\mu$ requires differentiation of the nonlinear function $K(u)$ present in (3) and, since the latter is usually available as a lookup table, it is approximated by an appropriate polynomial function of $u$. Fig. 5 illustrates a ninth-order polynomial approximation for $K(u)$, i.e.
$K(u)=\sum_{i=0}^{9}\alpha_{i}\,u^{i},$ (25)
for both an ideal actuator and the deteriorated (denoted real in Fig 5) actuator used later in the experimental validation.
Such approximation also ensures that for the case of the ideal valve $\frac{\partial K}{\partial u}>~0,\;\forall u\in[-1,1]$, as shown in Fig. 5. This condition will be useful later in proving the validity of the convergence conditions. The function of $\mu$ can now be selected as
@@
Moreover, since $\left|\displaystyle\frac{\partial\rho}{\partial u}\right|$ is bounded, there exists a Lipschitz constant $L_{u}>0$ such that
@@
Figure 5: Flow coefficient derivative for algorithm convergence condition and valve degradation effect.
Remark 3.
In the case of a non-ideal valve, e.g. due to wear, the term $\frac{\partial K}{\partial u}$ may take non-positive values as illustrated in Fig. 5, which may lead to violation of Condition 1. This will, in turn, result to a bounded input estimation error. An alternative yet more conservative choice is to select $\mu$ to be a positive constant.
@@ V Stability Analysis
The stability of the closed-loop system is analyzed by including the nonlinear controller, the dynamical mapping inversion, and the disturbance estimator. Let $\delta_{\underline{\mathbf{e}}}$ be the error associated with the calculation of $\ddot{e}$. The dynamics of the closed-loop tracking error $\underline{\mathbf{e}}$ is written as:
@@
[…]
$=$ $(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}})\,\underline{\mathbf{e}}+\boldsymbol{B_{FL}}\left(\boldsymbol{K_{FL}}\begin{bmatrix}0\\ -\frac{\tilde{F}_{d}}{M}\end{bmatrix}-\frac{\dot{\tilde{F}}_{d}}{M}\right)$
[…]
@@
[…]
Moreover, let $v^{*}=\rho(u^{*})$ be the required virtual controller and $u^{*}$ the related control input. Then, one can define
[…]
for some arbitrary $u$ and $v$. From (20) the dynamics of the input estimation error $\tilde{u}$ reads:
[…]
The entire closed-loop system can therefore be written as
$\underbrace{\begin{bmatrix}\dot{\underline{\mathbf{e}}}\\ \dot{\tilde{u}}\end{bmatrix}}_{\boldsymbol{\dot{\chi}}}=\underbrace{\begin{bmatrix}(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}})\,\underline{\mathbf{e}}+\boldsymbol{h_{FL}}(t,\tilde{u})\\ f_{2}(t,\underline{\mathbf{e}},\tilde{u})\end{bmatrix}}_{\boldsymbol{f_{\chi}}(t,\boldsymbol{\chi})}+\underbrace{\begin{bmatrix}\boldsymbol{I_{3}}\\ \boldsymbol{0_{1\times 3}}\end{bmatrix}}_{\boldsymbol{B_{\chi}}}\boldsymbol{\xi_{F_{d}}}.$ (34)
Lemma 1.
The unperturbed closed-loop system $\boldsymbol{\dot{\chi}}=\boldsymbol{f_{\chi}}(t,\boldsymbol{\chi})$ ($\boldsymbol{\xi_{F_{d}}}=\boldsymbol{0}$) is Uniformly Globally Asymptotically Stable (UGAS).
Proof.
The system can be seen as the feedback interconnection of the following subsystems:
[…]
as illustrated in Fig. 6.
The following statements are true:
  1. S1:
    The unperturbed subsystem $\Sigma_{1}$ $(\tilde{u}=0)$ is Uniformly GES (UGES).
  2. S2:
    The origin of the subsystem $\Sigma_{2}$ is UGES.
  3. S3:
    There exist two scalar, positive-definite functions $V_{FL},W$ such that
@@ 3. S3:
    $\frac{\partial V_{FL}(t,\underline{\mathbf{e}})}{\partial\underline{\mathbf{e}}}\boldsymbol{h_{FL}}(t,\tilde{u})=o(W(\underline{\mathbf{e}})),\;\forall(t,\underline{\mathbf{e}})\in[0,\infty)\times\mathbf{R}^{3}.$ (36)
The validity of S1 directly follows from the fact that $\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}}$ is Hurwitz by design and hence, with $\tilde{u}=0$, the system $\dot{\underline{\mathbf{e}}}=(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}})\,\underline{\mathbf{e}}$ has a UGES equilibrium at origin. The statement S2 is verified by the design of the estimator in (20), which under Conditions 1-3, guarantees exponential stability of the origin of the estimation error $\tilde{u}(t)$ at all times. Finally, consider the scalar, positive-definite Lyapunov function $V_{FL}=\underline{\mathbf{e}}^{T}P_{FL}\,\underline{\mathbf{e}}$, where $P_{FL}$ is the symmetric, positive-definite solution to the Lyapunov matrix equation
$(\boldsymbol{P_{FL}A_{FL}}-\boldsymbol{B_{FL}K_{fl}})+(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}})^{T}\boldsymbol{P_{FL}}=-\boldsymbol{I}$
associated to the UGES origin of the unperturbed $\Sigma_{1}$. Then
[…]
@@
$\leq\left\|2\boldsymbol{P_{FL}}\underline{\mathbf{e}}\right\|\,\left\|\boldsymbol{B_{FL}}\right\|\,L_{u}\,|\tilde{u}|\leq\underbrace{\left\|2\boldsymbol{P_{FL}}\right\|\,\left\|\boldsymbol{B_{FL}}\right\|\,L_{u}\,|\tilde{u}|_{\infty}}_{\eta>0}\,\left\|\underline{\mathbf{e}}\right\|$
$=\eta\|\underline{\mathbf{e}}\|^{2}\triangleq W(\|\underline{\mathbf{e}}\|).$ (37)
The last inequality implies $\frac{\partial V_{FL}(t,\underline{\mathbf{e}})}{\partial\underline{\mathbf{e}}}\boldsymbol{h_{FL}}(t,\tilde{u})=o(W(\underline{\mathbf{e}}))$, $\forall(t,\underline{\mathbf{e}})\in[0,\infty)\times\mathbf{R}^{3}$.
Following the approach introduced in [loriaFeedbackCascadeinterconnectedSystems2008], the unperturbed closed-loop system (35) can be seen as a cascaded interconnection of $\Sigma_{1}$ and $\Sigma_{2}$, where the solutions $\underline{\mathbf{e}}(t;\underline{\mathbf{e}}_{\boldsymbol{0}})$ of $\Sigma_{1}$ for all initial conditions $\underline{\mathbf{e}}_{\boldsymbol{0}}$ can be thought of as time-varying parameters of $\Sigma_{2}$. Then since statements S1, S2, S3 are true, by Proposition 1 in [loriaFeedbackCascadeinterconnectedSystems2008] the solutions of the unperturbed closed-loop system are globally UGAS. ∎
Proposition 1.
The solutions of the entire closed-loop system in Eq. (34) are Uniformly Ultimately Bounded (UUB).
Proof.
This follows directly from Lemma 1 and Lemma 9.3 in [khalilNonlinearSystems2002] since the bounded perturbation $\boldsymbol{\xi_{F_{d}}}$ is non-vanishing at the origin.
Figure 6: Closed loop system $\boldsymbol{\dot{\chi}}=\boldsymbol{f_{\chi}}(t,\boldsymbol{\chi})+\boldsymbol{B_{\chi}}\boldsymbol{\xi_{F_{d}}}$.
@@ VI Case Study
$J_{tot}\dot{v_{\beta}}=\overbrace{\left(A_{p}p_{p}-A_{r}p_{r}\right)r_{\beta}\sin\left(\psi\left(x_{c}\right)\right)}^{M_{h}}-M_{fric}-M_{ext}$ (38)
@@
where the equivalent mass $M$ now also accounts for blade inertia. (38) matches the dynamics of the cylinder dynamics presented in (7). The presence of a cyclic component in the variable distance that relates mechanical moment to force is minor, as shown in Fig. 7. Thus, the control design aims at tracking the angle while the cylinder dynamics represents a stable internal dynamics, under the assumption of having a stiff mechanical connection between cylinder and blade. Hence, there exists a static transformation relating $x_{c}$ to $\beta$ and with the pitch tracking error $e_{\beta}$ being the difference between pitch angle and a reference,
$e_{\beta}=\beta-r,$ (39)
the relative degree of the closed loop pitch control will be the same as for the closed loop cylinder piston position control loop analyzed above. With relative degree three of both, the third derivative of the new tracking error is needed,
$e_{\beta}^{(3)}=$ $-\frac{\dot{M}_{fric}+\dot{M}_{ext}}{J_{tot}}$ (40)
$-\frac{r_{\beta}A_{p}^{2}B}{J_{tot}V_{p}}v_{c}\sin{\psi_{\beta}(x_{c})}-\frac{r_{\beta}A_{r}^{2}B}{J_{tot}V_{r}}v_{c}\sin{\psi_{\beta}(x_{c})}$
$+\frac{r_{\beta}A_{p}B}{J_{tot}V_{p}}Q_{p}(u)\sin{\psi_{\beta}(x_{c})}+\frac{r_{\beta}A_{r}B}{J_{tot}V_{r}}Q_{r}(u)\sin{\psi_{\beta}(x_{c})}$
$+\underbrace{\frac{r_{\beta}}{J_{tot}}F_{h}\cos(\psi_{\beta}(x_{c}))\frac{d\psi}{dx_{c}}v_{c}}_{\tau(x)}-r^{(3)}\,\triangleq\,v.$ (41)
With respect to the general case presented in (8), two variations are present:
  • $1/M$ is replaced by $r_{\beta}\sin{\psi_{\beta}(x_{c})}/J_{tot}$.
@@
However, being $\psi_{\beta}(x_{c})$ is approximately constant, the following assumptions are made:
  • $\psi_{\beta}(x_{c})\approx\textit{const}$
  • $\frac{d\psi_{\beta}}{dx_{c}}\approx 0\rightarrow\tau(x)\approx 0$.
VI-A Simulation Setup
The designed control actions are tuned on a high-fidelity simulator, which includes the general dynamics presented in (1) and (38), and also encompass the control valve dynamics. The measured signals are subjected to noise, simulated to match the levels in the test rig. The velocities ($v_{c}$ and $v_{\beta}$) are obtained via numerical differentiation since they are not usually available as sensor signals in a real plant. The reference for the pitch angle and the wind load $M_{ext}$ are obtained from a wind turbine high-fidelity simulator based on OpenFAST, widely used in wind turbine research and development [Jonkman2013].
VI-B Experimental setup
The proposed algorithm is further tested on a full scale pitch system from a real wind turbine with a rated power of 3 MW. A picture of the setup is presented in Fig. 8, and further details can be found in [pedersenOnlineLoadLeakage2025].
Figure 8: Full-scale experimental setup from a wind turbine: $104\,\mathrm{kg}$ hydraulic cylinder with a rated force of $150\,\mathrm{kN}$, blade mount and pitch bearing with an inertia of $280\,\mathrm{kg\,m}^{2}$. The 80 m blade is not attached, but its inertia is emulated through the action of a load cylinder, which excites the setup in real time, and also includes the OpenFAST calculated blade moment.
For practical reasons, the test rig does not include the wind turbine blade. However, an additional load cylinder is available to both replicate the wind load from OpenFAST and digitally reproduce the blade’s inertia. Fig. 9 presents a detailed diagram of the test rig’s composition.
In the real turbine, the total inertia is given by the inertia of the pitch test rig ($J_{tr}$) plus the inertia of the blade ($J_{b}$). One can write:
$J_{tr}\dot{v_{\beta}}$ $=M_{h}-M_{fric}-M_{ext}-J_{b}\dot{v_{\beta}}$ (42)
$\approx M_{h}-M_{fric}-\underbrace{\left(M_{ext}+J_{b}\,\ddot{r}\right)}_{M_{h,l}},$
by assuming $r\approx\beta$, and avoiding double differentiating the noisy position measurement. Thus, the inertia of the blade is reproduced by injecting an additional mechanical torque by means of the load cylinder.
The real-time implementation of the control action is performed using a Speedgoat real-time module for MATLAB/Simulink.
VI-C Controllers Design
The control action for the baseline feedback linearization controller and for the proposed algorithm are derived by adapting the methods described in Sec. III and Sec. IV, respectively.
VI-C1 Standard controller